OpenStudy (anonymous):

multiply and simplify the expression (3w-4)(4w-6)

OpenStudy (anonymous):

@Jhannybean @jim_thompson5910 @tkhunny @Zale101

OpenStudy (tkhunny):

Have you considered multiplying? Please show your work.

OpenStudy (anonymous):

i was thinking 12w-24 but im not sure

OpenStudy (tkhunny):

Can't be. There must be a \(w^{2}\) in there, somewhere. Give it another go.

OpenStudy (anonymous):

\[12w ^{2}-24\]

OpenStudy (jhannybean):

\[\large (3w-4)(4w-6)\]\[\large \color{green}{3w}(4w-6) -\color{green}4(4w-6)\]\[\large (3\cdot 4)(w\cdot w) - (3w \cdot 6) -(4 \cdot 4w) -4(-6)\] Simplify and reduce.

OpenStudy (anonymous):

\[12w ^{2}-34w+24\]